A gas mixture contains 12.2 g of N2,5.00 g of H2, and 7.10 g of CH4. What is the mole fraction of CH, in the mixture?
Answer #1

The molecular weights of N2, H and CH4 are, 28, 2 and 16 g/mol, respectively.

Let us now calculate the number moles of each species by using following formula

Number of moles = Weight in g / Molecular weight g/mol

Number of moles of N2 = 12.2 g / 28 g/mol = 0.435 moles

Number of moles of H2 = 5 / 2 = 2.5 moles

Number of moles of CH4 = 7.10 g / 16 = 0.443 moles

The total number of moles present in the mixture = 0.435 + 2.5 + 0.443 = 3.38 moles

The mole fraction for CH4 = number of moles of CH4 / Total number of moles

= 0.443 / 3.38

= 0.131

Therefore, mole fraction of CH4 in the given mixture is, 0.131.